OpenStudy (anonymous):

@iGreen @akonkel @Adi3 i need help really bad i will meadel and fan choose the graph that shows the solution of the inquality on the number line c<-1

OpenStudy (anonymous):

hold on i have to get the graphs

OpenStudy (igreen):

Open dot will be for < or > Closed dot will be for \(\le\) or \(\ge\)

OpenStudy (igreen):

c < -1 The line will start on '-1'

OpenStudy (igreen):

\(<\) and \(\le\) will go to the left \(>\) and \(\ge\) will go to the right.

OpenStudy (igreen):

Which do you think is correct?

OpenStudy (adi3):

wht do u think

OpenStudy (anonymous):

yhe right

OpenStudy (anonymous):

i think it is going to the right

OpenStudy (anonymous):

any body there i am going to give you the rest of the graphs

OpenStudy (igreen):

Open dot will be for < or > Closed dot will be for \(\le\) or \(\ge\) \(<\) and \(\le\) will go to the left \(>\) and \(\ge\) will go to the right. Our inequality is \(c < -1\)

OpenStudy (igreen):

So, will it have an open dot or a closed dot?

OpenStudy (anonymous):

closed

OpenStudy (adi3):

I will try to explain bc net is really slow bc its raining heavily here

OpenStudy (igreen):

@malikjipock Is this it? \(c \le -1\) or is this it? \(c < -1\)

OpenStudy (adi3):

noooooooooooooooo

OpenStudy (anonymous):

c<-1

OpenStudy (anonymous):

i green are u there here it is

OpenStudy (igreen):

Okay, so is it going to be an open dot or a closed dot?

OpenStudy (anonymous):

cloesed

OpenStudy (igreen):

No.. Open dot will be for < or > Closed dot will be for \(\le\) or \(\ge\) Since it's '<' it will be an open dot..

OpenStudy (igreen):

Will it go to the left or to the right?

OpenStudy (anonymous):

left

OpenStudy (igreen):

Yes, so our line will have an OPEN DOT and it will be going to the LEFT, so which graph should be correct?